assume that the salaries of the elementary school teachers in the united states are normally distributed with a mean of $31,450 and a standard deviation of $2,995. What is the cutoff salary for teachers in the bottom 15%?


1
Expert's answer
2021-05-07T14:17:45-0400

We have given,

"\mu = 31450"

"\sigma = 2995"

z score = "-1.036"

We use the formula,


"z = \dfrac{X - \mu}{\sigma}"


"-1.306 = \dfrac{X-31450}{2995}"


"-3911.47 = X - 31450"


"X = 27538.53"

Hence, the cut off salary "= 27538.53$"